1 THIS IS THE FIRST SECTION 1

1 This is the First Section

This is the first paragraph. It is indented, since I’m using \usepack-

age{indentfirst} in my LYX preamble.

• some item

• another item

• yet another item Surprise, this paragraph is NOT indented.